The risk-free rate of return is 3.68 percent and the market risk premium is 7.84 percent. What is the expected rate of return on a stock with a beta of 1.32?
Answer #1

Rf=0.0368

Rm-Rf=0.0784

Beta(B)=1.32

Expected return(E)=Rf+B*(Rm-Rf)=0.0368 +(Rm-Rf)*B=0.140288 i.e 14%